OpenStudy (lena772):

David's company reimburses his expenses on food, lodging, and conveyance during business trips. The company pays $60 a day for food and lodging and $0.65 for each mile traveled. David drove 600 miles and was reimbursed $3390. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did David spend on this trip? (1 points)

OpenStudy (ranga):

Let x be the number of days on the trip. The company pays $60 a day for food and lodging. Therefore for x days the company would pay: 60 * x The company pays $0.65 for each mile traveled. For traveling 600 miles the company would have paid: 600 * 0.65 = $390 Total expenses = 60x + 390 = 3390 is your equation.

OpenStudy (lena772):

Ok i know how to solve but what is the property of equality

OpenStudy (lena772):

@ranga

OpenStudy (ranga):

I think property of equality refers to adding or subtracting same things to both sides of the equality keeps the equation in tact. Same with multiplication and division.

OpenStudy (lena772):

Oh ok ik how to do that

OpenStudy (ranga):

Property of equality: 1. Addition: If a = b then a + c = b + c. 2. Subtraction: If a = b then a – c = b– c. 3. Multiplication: If a = b then ac = bc. 4. Division: If a = b and c ≠ 0 then a/c = b/c.

OpenStudy (ranga):

alright.

OpenStudy (lena772):

60x+390=3390 60x+390-390=3390-390 60x=3000 60x/60=3000/60 x=50

OpenStudy (lena772):

Is that right?

OpenStudy (lena772):

@ranga

OpenStudy (ranga):

Yes.

OpenStudy (lena772):

do i have to put the name of each quality as i go?

OpenStudy (ranga):

Looks like they are asking for justification in each step.

OpenStudy (lena772):

How do I justify it?

OpenStudy (ranga):

Look in your text/notes how they do it and follow the same procedure to get full credit. It may be something like: 60x+390-390=3390-390 by subtraction property of equality

OpenStudy (lena772):

Thank you!

OpenStudy (lena772):

So David spent 50 days on the trip?

OpenStudy (ranga):

Yes. 50 days.

OpenStudy (anonymous):

are you doing flvs? @Lenna772

OpenStudy (lena772):

yup

OpenStudy (anonymous):

what lesson thingy are you on